OpenStudy (anonymous):

A math club is researching a golf tournament fund-raiser. It will cost $1,000 to host the tournament. If it rains, the club will lose the investment. If it is sunny, it is expected that the club will collect $4,500 from the participants. If the chance of rain is 20%, what is the expected value for the tournament? @DanJS

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

Expected value: \[ E[X] =\mu =\sum_{i} x_i\cdot \Pr(X=x_i) \]

OpenStudy (anonymous):

In this case, we have two \(i\) values. \(i=1\) is the case that it rains. \(i=2\) is the case that it doesn't rain.

OpenStudy (anonymous):

Here are the earnings:\[ x_1= -$1000\\ x_2 = $4500-$1000 = $3500 \]

OpenStudy (anonymous):

Here are the probabilities: \[ \Pr(X=x_1) = 0.20\\ \Pr(X=x_2) = 1-0.20 = 0.80 \]

OpenStudy (anonymous):

Therefore... \[ E[X] = 0.20\cdot (-$1000) + 0.80\cdot ($3500) \]

OpenStudy (anonymous):

2600 and 3400 are the only answers that are positive

OpenStudy (anonymous):

All that remains is to multiply and add.

OpenStudy (anonymous):

2600 thank you!

OpenStudy (anonymous):

But do you understand how I did it...

OpenStudy (anonymous):

I remembered the formula, and that you took the two probabilities of the occurrences and multiplied them